OpenStudy (anonymous):

I am working on averaqe rate of change of a function and i need help simplifying this [sin 3pi/2 -sin5pi/6] / 3pi/2 -5pi/6

OpenStudy (turingtest):

\[\Large{\sin(\frac{3\pi}2)-\sin(\frac{5\pi}2)\over\frac{3\pi}2-\frac{5\pi}2}\]

OpenStudy (turingtest):

right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I would evaluate sine functions first.

OpenStudy (anonymous):

wasn't that a 5pi/6?

OpenStudy (anonymous):

oh yeah. i missed that. thanks

OpenStudy (anonymous):

its 5pi/6

OpenStudy (anonymous):

k... what do you get for sin(3pi/2) and sin(5pi/2)?

OpenStudy (anonymous):

You can use unit circle, reference triangles, or even a calculator in radian mode.

OpenStudy (anonymous):

the choice of ansers given were not completely simplified and thats whats confusing. When i use a calcultor, it simplifies it to 88.

OpenStudy (anonymous):

Oh, I was just talking about the sines for the calculator.

OpenStudy (anonymous):

but the choice of answers are all in fractions with pi in them

OpenStudy (anonymous):

sin(3pi/2)=-1 What is sin(5pi/2)?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Yes, so you have -1-1 in the numerator... which simplifies to?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Yes, now find the lowest common denominator of the fractions in the denominator.

OpenStudy (anonymous):

Both 2 and 6 go into what number?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Yes, so change 3pi/2 into sixths by multiplying top and bottom by the number that makes 2 into 6...which is 3.

OpenStudy (anonymous):

\[\frac{3}{3}\frac{3\pi}{2}=\frac{9\pi}{6}\]

OpenStudy (anonymous):

9pi/6

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yes, now combine the fractions below by subtracting their numberators.

OpenStudy (anonymous):

15pi/18

OpenStudy (anonymous):

-6pi

OpenStudy (anonymous):

\[\frac{9\pi}{6}-\frac{5\pi}{6}=\frac{4\pi}{6}\]

OpenStudy (anonymous):

Here is what I have all together now.\[\frac{-2}{\frac{4\pi}{6}}\]

OpenStudy (anonymous):

To divide fraction, invert the denominator and multiply by the numerator.\[-2(\frac{6}{4\pi})\]and reduce.

OpenStudy (anonymous):

Any questions, or maybe a possible answer?

OpenStudy (anonymous):

when you multipied the denom and numer of the first fraction by 3, why didnt you do it to the second fraction?

OpenStudy (anonymous):

I was changing its form so that I could subtract the fractions. A common denominator is required to subtract fractions. By multiplying the same number top and bottom 3/3, it was like multiplying by one. It changes its form without changing its value.

OpenStudy (anonymous):

ok i got it. because 9pi/6 is the same as 3pi/2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and at that point it is compatable with 5pi/6... so we can subtract them.

OpenStudy (anonymous):

ok. their final answer came out to be -9/4pi

OpenStudy (anonymous):

bummer, that is not what I got... let me double check.

OpenStudy (anonymous):

Oh I see... I missed that your original sine was 5pi/6 as well. I thought it was 5pi/2.

OpenStudy (anonymous):

oh yeah, your right

OpenStudy (anonymous):

sin(5pi/6)=1/2

OpenStudy (anonymous):

-1-1/2 in the numerator gives -3/2

OpenStudy (anonymous):

\[\frac{-\frac{3}{2}}\frac{4pi}{6}\]

OpenStudy (anonymous):

editor is giving me problems... -3/2 in the numerator 4pi/6 in the denominator

OpenStudy (anonymous):

Invert and multiply denominator...\[-\frac{3}{2}\frac{6}{4\pi}\]

OpenStudy (anonymous):

Bummer, it still looks bad on my screen. But yes, in the end it comes down to -9/4pi.

OpenStudy (anonymous):

yes it sure does . thanks a lot

OpenStudy (anonymous):

No problem. It was fun until I started getting glitches.

OpenStudy (anonymous):

lol. i know. its difficult typing those characters in

OpenStudy (anonymous):

i got another similar problem